OpenStudy (anonymous):

how do i do: 4(3m +2)??? I DO NOT want just an answer, I need someone to actually show me how to solve it.

OpenStudy (anonymous):

Distriubute the 4 into both terms of the binomial... 4(3m + 2) = 4(3m) + 4(2) Multiply these out and you're done.

OpenStudy (anonymous):

see here : use distributive property which states that x(a+b) = xa + xb here: 4(3m+2) = 4(3m) + 4(2) 4(3m+2) = 12m + 8 i hope u would have understood this